# Turning Down the Heat: A Critical Analysis of Min-p Sampling in Language Models

### v1

@@ 3.1 Thorough Hyperparameter Sweep on GSM8K Contradicts Claim of Min-p’s Superiority
To test whether min-p indeed achieves superior performance, we conducted an extensive analysis on GSM8K, sweeping the following models, samplers, hyperparameters and sampling seeds:
@@
Due to our compute budget, we only evaluated GSM8K (albeit under two prompt formats, for reasons explained below). This sweep and the sweep below required $\sim 6000$ Nvidia A100-hours. GSM8K contains a subset of samples with ambiguous language or incorrect labels that have since been identified and cleaned (Vendrow et al., 2025) and that models may have been trained on GSM8K (Zhang et al., 2024), but we used GSM8K nonetheless for consistency with the original paper. We similarly used EleutherAI’s LM Eval Harness (Gao et al., 2021; Biderman et al., 2024).
@@
After we showed these results to the authors, they informed us that we had run our experiments using the “Llama" formatting of GSM8K prompts as we used the command from the authors’ public Colab notebook; the authors clarified that "Llama" formatting should be used only for Llama models. We then reran our experiments using standard formatting of GSM8K prompts. The results were nearly identical (Appendix B), with one small difference: min-p does produce higher scores for 2 of 12 language models. Again, we conclude min-p does not outperform other samplers on either formatting of GSM8K when controlling for hyperparameter volume.
@@ Appendix A Examples of Human Qualitative Responses Favoring Basic Sampling Over Min-P Sampling
In Section 2.3, we described how qualitative responses from many human participants in the original paper’s study favored basic sampling. Direct quotes from human evaluators favoring basic sampling are provided below. In the study, basic sampling was called “Model A”; for clarity, we substituted below for clarity):

### v2

# Min-p, Max Exaggeration: A Critical Analysis of Min-p Sampling in Language Models

@@ 3.1 Thorough Hyperparameter Sweep on GSM8K Contradicts Claim of Min-p’s Superiority
To test whether min-p indeed achieves superior performance, we conducted an extensive analysis on GSM8K CoT, sweeping the following models, samplers, hyperparameters and sampling seeds:
@@
Due to our compute budget, we only evaluated GSM8K CoT (albeit under two prompt formats, for reasons explained below). This sweep and the sweep below required $\sim 6000$ Nvidia A100-hours. GSM8K contains a subset of samples with ambiguous language or incorrect labels that have since been identified and cleaned (Vendrow et al., 2025) and that models may have been trained on GSM8K (Zhang et al., 2024), but we used GSM8K nonetheless for consistency with the original paper. We similarly used EleutherAI’s LM Eval Harness (Gao et al., 2021; Biderman et al., 2024).
@@
After we showed these results to the authors, they informed us that we had run our experiments using the “Llama" formatting of GSM8K CoT prompts as we used the command from the authors’ public Colab notebook; the authors clarified that "Llama" formatting should be used only for Llama models. We then reran our experiments using standard formatting of GSM8K CoT prompts. The results were nearly identical (Appendix B), with one small difference: min-p does produce higher scores for 2 of 12 language models. Again, we conclude min-p does not outperform other samplers on either formatting of GSM8K CoT when controlling for hyperparameter volume.
@@ Appendix A Examples of Human Qualitative Responses Favoring Basic Sampling Over Min-P Sampling
In Section 2.3, we described how qualitative responses from many human participants in the original paper’s study favored basic sampling. Direct quotes from human evaluators favoring basic sampling are provided below. In the study, basic sampling was called “Model A”; for clarity, we substituted the pseudonyms for the actual sampling methods):